If 318 g of Iron (.444 J/g °C) at 118°C is placed in 218 g of water at 17°C in an insulated container and both are allowed to come to same temperature. What will the temperature be?
Answer #1

Ans :

The heat lost by the iron will be gained by water

Heat = mass x specific heat x change in temperature

specific heat water = 4.186 J/goC

Let the final temperature be T

putting values :

= 318 x 0.444 x (118 - T) = 218 x 4.186 x (T - 17)

= 16660.656 - 141.192 T = 912.548 T - 15513.316

= 16660.656 + 15513.316 = 912.548 T + 141.192 T

= 32173.972 = 1053.74 T

T = 30.53oC

So the temperature will be 30.53oC